### 24 Mar 2024 v5

# ALI-DPFL: Differentially Private Federated Learning with Adaptive Local Iterations

Abstract
Federated Learning (FL) is a distributed machine learning technique that allows model training among multiple devices or organizations by sharing training parameters instead of raw data. However, adversaries can still infer individual information through inference attacks (e.g. differential attacks) on these training parameters. As a result, Differential Privacy (DP) has been widely used in FL to prevent such attacks.
We consider differentially private federated learning in a resource-constrained scenario, where both privacy budget and communication rounds are constrained. By theoretically analyzing the convergence, we can find the optimal number of local Differential Privacy Stochastic Gradient Descent (DPSGD) iterations for clients between any two sequential global updates. Based on this, we design an algorithm of Differentially Private Federated Learning with Adaptive Local Iterations (ALI-DPFL). We experiment our algorithm on the MNIST, FashionMNIST and Cifar10 datasets, and demonstrate significantly better performances than previous work in the resource-constraint scenario. Code is available at https://github.com/KnightWan/ALI-DPFL.
Index Terms:
differential privacy, federated learning, adaptive, convergence analysis, resource constrained
I Introduction
In federated learning, each client normally downloads the global model from the center sever, performs local iterations, and uploads the resulted training parameters back to the center sever. The center sever then updates the global model accordingly. The above steps repeat until the global model converges [mcmahan2017communication]. In this way, the global model reaches convergence by communicating only training parameters instead of raw data. Nevertheless, some studies have shown that federated learning still carries privacy risks. Training parameters, such as gradient values, can be used to recover a portion of the original data [zhu2019deep] or infer whether specific content originates from certain data contributors [song2017machine]. Melis et al. [LucaMelis2022ExploitingUF] also demonstrated that participants’ training data could be leaked by shared models. Therefore, additional measures need to be taken to protect data privacy in federated learning.
Many works[AnttiKoskela2018LearningRA, LeTrieuPhong2017PrivacyPreservingDL, jayaraman2019evaluating, LiyaoXiang2019DifferentiallyPrivateDL] have demonstrated that the technique of differential privacy (DP) could protect machine learning models from unintentional information leakage. In differentially private federated learning (DPFL), each client executes a certain number of local stochastic gradient descent with differential privacy (DPSGD) iterations before performing a global aggregation. For each global aggregation, DPFL consumes privacy budget according to the number of local iterations and communicates one round of training parameters, and thus DPFL should be constrained by both the privacy budget and communication rounds. However, current DPFL schemes typically optimize the global model only with the constraint of a limited privacy budget, while overlooking the constraint of communication rounds.
(a) Iters=150
(b) Iters=300
(a) Iters=150
(b) Iters=300
(c) Iters=600
[…]
Figure 2: The Framework of Differentially Private Federated Learning with Adaptive Local Iterations
Existing DPFL schemes usually choose empirically a fixed number of local iterations $\tau$ for each global update, usually, $\tau=1$. They treat communication rounds as unlimited, but in practice, the number of communication rounds are often constrained by limited resources such as time and bandwidth. As shown in Fig. 1, the optimal fixed $\tau$ value corresponding to the highest accuracy is not a fixed value across different datasets and iterations. It is also indicated that a fixed $\tau$ may fail to achieve satisfactory convergence performances [LiT2020FedProx, WangJ2020FedNova, WangS2019ResourceConstrained]. Therefore, for the first time, in this paper we focus on finding a DPFL scheme with adaptive local iterations to achieve good performances when both privacy budget and communication rounds are constrained.
Through the convergence analysis, we derive a convergence bound related to the number $\tau$ of local iterations. Then, we propose a differentially private federated learning scheme with adaptive local iterations (ALI-DPFL), as shown in Fig. 2, which finds the optimal value ($\tau^{*}$) in each communication round that minimizes the convergence bound. This value is then used as the number of local iterations in the next round, enabling the algorithm to converge faster under the constraints of privacy budget and communication rounds. The main contributions of this paper are as follows:
  1. 1.
    We analyze DPFL convergence under constrained privacy budgets and communication rounds, presenting a novel convergence bound related to the number of local iterations.
  2. 2.
    Leveraging this bound, we introduce the ALI-DPFL algorithm, employing adaptive local iterations for high accuracy and fast convergence in resource-constrained settings.
  3. 3.
    Our extensive experiments showcase our approach’s superiority over three leading DPFL algorithms across IID and Non-IID data distributions..
The rest of the paper is organized as follows. Section II reviews prior research. Section III introduces the preliminary knowledge. Section IV proposes the ALI-DPFL algorithm through analyzing convergence bound of the DPFL algorithm and proves our algorithm satisfy differential privacy. The experimental results are presented in Section LABEL:sec:experiment, followed the conclusion in Section LABEL:sec:conclusion.
II Related Work
II-A DPFL
In previous studies, two different types of differential privacy have been incorporated into federated learning to protect privacy: client-level differential privacy (CL-DP) and sample-level differential privacy (SL-DP).
CL-DP. In the CL-DP setting, the model’s performance remains stable despite the addition or removal of one client during training [HBrendanMcMahan2017LearningDP]. Enayat et al. [Enayat2023PrivateFLwithAutotunedCompression] utilized CL-DP to dynamically adjust the compression rate based on training loss, thereby reducing communication rounds in federated learning. Zhou et al. [zhou2023exploring] explored the impact of local iterations when employing the Laplace mechanism for perturbation.
SL-DP. In the SL-DP setting, each client possesses a set of data samples. Importantly, the addition or removal of a data sample during training has minimal impact on the model’s performance [MartnAbadi2016DeepLW, XiWu2016BoltonDP]. Fu et al. [FuJ2022AdapDPFL] introduced adaptive gradient clipping, tailoring it for different clients and communication rounds based on observed changes in loss values. Raouf et al. [raouf2021ConstrainedDPFLforlowBandwidth] leveraged SL-DP to enhance model accuracy by updating only a fixed subset of model weights while leaving others unchanged, thereby reducing bandwidth consumption for both upstream and downstream communication. Kang Wei et al. [KangWei2021UserLevelPF] proposed the adaptive privacy budget allocation. Zhou et al. [zhou2021optimizing] provided the optimal value for the total local iterations through convergence analysis and analyzed the impact of the total number of iterations on convergence performance and privacy protection. Our work focuses on the SL-DP setting, which is the most popular in the real world.
II-B Adaptive local iterations in FL
There are currently numerous adaptive federated learning algorithms aimed at improving model performance in heterogeneous data settings. For example adaptive model aggregation [Sunwoo2023LayerWise, jayaram2022adaptive], adaptive optimizers [reddi2020FedOPT], adaptive Network Layering [karimi2021layer] and adaptive client sampling [Boxin2023AddressingBudget], etc. Wang et al[WangS2019ResourceConstrained] improved the federated learning algorithm by considering scenarios where clients and the server consume the same type of resources. They achieved faster convergence of the global model under resource constraints by dynamically adjusting local iterations.
Some studies like [WangS2019ResourceConstrained] addressed optimal iteration rounds in FL under limited resources but overlooked differential privacy. Others, such as [zhou2023exploring][zhou2021optimizing], calculated optimal iteration rounds in DPFL but neglected communication round constraints. This paper, to our knowledge, represents the first attempt to utilize SL-DP convergence analysis to dynamically determine the number of local iterations, achieving rapid convergence.
III Preliminaries
In the beginning, we list some notations used in this paper in Table I. Symbols that have not been mentioned before will be defined later in the following sections.
TABLE I: Summary of main notations
$F(\mathbf{w}$)
Global loss function
$F_{i}(\mathbf{w}$)
Local loss function for client $i$
$t$
Iteration index
$k$
Global aggregation index
$\mathbf{w}^{i}_{t}$
Local model parameter at node $i$ at iteration $t$
$\mathbf{w}_{t}$
Global model parameter at iteration $t$
$\mathbf{w}^{*}$
True optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$
Gradient descent step size
$\tau$
Number of local update iterations between two global aggregations
$\tau^{*}$
Dynamic optimal $\tau$ in a single round
$T$
Total number of local iterations
$K$
Total number of global aggregation, equal to $T/\tau$
$p_{i}$
weight of client $i$
$R_{s}$
Maximum communication rounds
$R_{c}$
Maximum total local iterations
$L$
Lipschitz parameter
$\mu$
Strongly convex parameter
$\sigma$
Standard deviation of Gaussian distribution
$d$
The dimension of model
$C$
Clipping bound
$q$
Sampling rate of batch
|||||$\cdot$||| | ⋅ | |
$L_{2}$-norm
III-A Differential privacy
Differential privacy is a rigorous mathematical framework that formally defines the privacy loss of data analysis algorithms. Informally, it requires that any changes to a single data point in the training dataset can not cause statistically significant changes in the algorithm’s output.
Definition 1
[…]
$\operatorname{Pr}\left[\mathcal{M}\left(\mathcal{D}_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(\mathcal{D}_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta$ (1)
III-B Differentially private stochastic gradient descent
Differentially Private Stochastic Gradient Descent (DPSGD) stands as a widely embraced deep learning algorithm offering privacy guarantees[MartnAbadi2016DeepLW, MingyiHong2020UnderstandingGC, ZhiqiBu2022Automatic, XiaodongYang2022Normalized, TianyuXia2023DPLwithPAC]. Specifically, at each iteration $t$, a batch of data $\mathcal{B}_{t}$ is sampled from the dataset using a sampling rate $q$. Subsequently, for each batch, $s$ is a sample thereof. Following the clipping of per-sample gradients, Gaussian noise with a mean of 0 is incorporated into the gradients in the subsequent step:
$\bar{g}_{t}(s)$ $=\textbf{clip}\left(\nabla F(\mathbf{w}_{t},s)\right)$
$=\nabla F(\mathbf{w}_{t},s)/\max\left(1,\frac{\|\nabla F(\mathbf{w}_{t},s)\|}{C}\right)$ (2)
$\tilde{g}_{t}=\frac{1}{|\mathcal{B}_{t}|}\left(\sum_{s\in\mathcal{B}_{t}}\bar{g}_{t}(s)+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I}_{d})\right)$ (3)
Here, $\sigma$ represents the noise multiplier, while $C$ denotes the clipping bound. Subsequently, gradient descent is executed based on the batch-averaged gradients. The initial models being randomly generated and independent of the sample data, combined with the batch-averaged gradients adhering to differential privacy, ensures that the resulting models also adhere to differential privacy, attributable to the post-processing property[jiefu2023DPSUR].
III-C Federated Learning with Differential Privacy
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier $\sigma$, local iterations $\tau$, clipping bound $C$, sampling rate $q$
Output: the final trained model $\mathbf{w}(K)$
1 Initialize $k=0$, $\mathbf{w}_{0}=\text{Initial()}$();
2 while $k<K$ do
3       for $i\in[1,2,\cdots,N]$ parallel do
4             for $j\in[1,2,\cdots,\tau]$ do
5                   $t=(k-1)\tau+j$;
6                   Sample randomly a batch $\mathcal{B}^{i}_{t}$ with sampling rate $q$;
7                   for $s\in\mathcal{B}^{i}_{t}$  do
8                         $g^{i}_{t}(s)\leftarrow\nabla F_{i}(\mathbf{w}^{i}_{t},s)$;
9                         $\bar{g}^{i}_{t}(s)\leftarrow g^{i}_{t}(s)/\max(1,\frac{\|g^{i}_{t}(s)\|}{C})$;
10
11                  $\widetilde{g}^{i}_{t}\leftarrow\sum\limits_{s\in{\mathcal{B}^{i}_{t}}}\bar{g}^{i}_{t}(s)+\mathcal{N}(0,\sigma^{2}{C}^{2}\mathbf{I}_{d})$;
12                   $\mathbf{w}^{i}_{t+1}=\mathbf{w}^{i}_{t}-\eta\cdot\frac{\widetilde{g}^{i}_{t}}{|\mathcal{B}^{i}_{t}|}$;
13
14      Receive $\text{Receive }\mathbf{w}^{i}_{t+1}$;
15       $k=k+1$;
16       $\mathbf{w}(k)=\sum_{i=1}^{N}p_{i}\mathbf{w}^{i}_{t+1}$;
17       Broadcast $\text{Broadcast }\mathbf{w}(k)\text{ to each client}$ ;
18
return $\mathbf{w}(K)$
Algorithm 1 Federated Learning with Differential Privacy
In a typical federated learning system, a server and $N$ clients form the fundamental components[mcmahan2017communication][mcmahan2016federated]. Each client, denoted as $\mathcal{C}_{i}$, maintains a local database $\mathcal{D}_{i}$, where $i\in\{1,2,\dots,N\}$. The quantity $|\mathcal{D}_{i}|$ signifies the size of the dataset at client $i$, and the total dataset size is denoted as $|\mathcal{D}|=\sum_{i=1}^{N}|\mathcal{D}_{i}|$. Let $p_{i}\triangleq\frac{|\mathcal{D}_{i}|}{|\mathcal{D}|}$ represent the weight of client $i$. The server’s objective is to learn a model using data distributed across the $N$ clients. Each active client engaged in local training endeavors to discover a vector $\mathbf{w}$ that represents the neural network model minimizing a specific loss function. Subsequently, the server aggregates the weights from the $N$ clients, a process that can be expressed as:
$\mathbf{w}(k)=\sum_{i=1}^{N}p_{i}\mathbf{w}^{i}_{t}$ (4)
Here, $\mathbf{w}^{i}_{t}$ signifies the parameter vector acquired from client $i$ at iteration $t$, while $\mathbf{w}(k)$ denotes the server-aggregated parameter vector at communication round $k$. The objective is to minimize $F(\mathbf{w})$, in other words, to seek:
$\mathbf{w}^{*}\triangleq\arg\min F(\mathbf{w})$ (5)
The function $F(\cdot)$ is defined as follows:
$F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w})$ (6)
Here, $F_{i}(\cdot)$ represents the local loss function of client $i$.
As depicted in Algorithm 1, the differentiating factor between the DPFL algorithm and traditional federated learning algorithms resides in the utilization of DPSGD instead of SGD. Specifically, at line 1, the DPFL algorithm executes per-sample clipping, and at line 1, it introduces noise to the clipped gradients.
IV ALI-DPFL Algorithm
In this section, we introduce our ALI-DPFL algorithm, which is based on the insights gleaned from the convergence analysis. In Section IV-A, we conducted a convergence analysis of the traditional DPFL algorithm to investigate the relationship between the convergence upper bound and $\tau$, resulting in Equation (LABEL:eq:ca). Leveraging the outcomes of the convergence analysis, we derived Lemma LABEL:lem:tau_0 and Lemma LABEL:lem:tau_star. Subsequently, in Section LABEL:subsec:overview, we outline the workflow of ALI-DPFL. Furthermore, in Section LABEL:subsec:algo_alidpfl, we constructed Algorithm LABEL:alg:ALI-DPFL based on Lemma LABEL:lem:tau_0 and Lemma LABEL:lem:tau_star. Finally, in Section LABEL:subsec:privacy_analysis, we demonstrated that Algorithm LABEL:alg:ALI-DPFL satisfies the conditions for differential privacy.
IV-A Motivation: Convergence Analysis of DPFL
To investigate the impact of $\tau$ on the convergence rate, we delve into the convergence of Algorithm 1 and draw insights on selecting the optimal local iteration count $\tau^{*}$ based on the convergence upper bound. Prior to substantiating this, we establish several assumptions:
Assumption 1
$F_{1},\cdots,F_{N}$ are all L-smooth[watson1964smooth]: for all $\mathbf{v}$ and $\mathbf{w},F_{i}(\mathbf{v})\leq F_{i}(\mathbf{w})+(\mathbf{v}-\mathbf{w})^{T}\nabla F_{i}(\mathbf{w})+\frac{L}{2}\|\mathbf{v}-\mathbf{w}\|^{2}$.